A group of market women sell at least one of yam, plantain and maize. 12 of them sell maize, 10 sell yam and 14 sell plantain. 5 sell plantain and maize, 4 sell yam and maize, 2 sell yam and plantain only while 3 sell all the three items. How many women are in the group?
Answer by ikleyn(52750)

.
A group of market women sell at least one of yam, plantain and maize.
12 of them sell maize, 10 sell yam and 14 sell plantain.
5 sell plantain and maize, 4 sell yam and maize, 2 sell yam and plantain only
while 3 sell all the three items. How many women are in the group?
~~~~~~~~~~~~~~~

We are given a set of women, which consists of 3 intersecting subsets, 

    - Y (sell yam and may be something else),
    - P (sell plantain and may be something else),
    - M (sell maize and may be something else).


About these sets and their intersections, we are given the following information

    n(M) = 12,                                                 (1)
    n(Y) = 10,                                                 (2)
    n(P) = 14,                                                 (3)

    n(PM) = 5       (for intersection P and M)                 (4)
    n(YM) = 4       (for intersection Y and M)                 (5)
    n(YP\YMP) = 2   ((for intersection Y and P) minus YMP)     (6)

    n(MYP) = 3      (for triple intersection Y, M and P).      (7)


Now, from (6) and (7), we get  

    n(YP) = n(YP\YMP) + n(YMP) = 2 + 3 = 5.                    (8)


    +-------------------------------------------+
    |    They want we determine n(M U Y U P).   |
    +-------------------------------------------+


Use the Inclusive-Exclusive principle formula

    n(M U Y U P) = n(M) + N(Y) + n(P) - n(PM) - n(YM) - n(YP) + n(YMP).


Substitute here all known quantities and get the answer

    n(M U Y U P) = 12 + 10 + 14 - 5 - 4 - 5 + 3 = 25.


ANSWER.  There are 25 women in the group.

Solved.

------------------

On inclusion-exclusion principle,  see this Wikipedia article

https://en.wikipedia.org/wiki/Inclusion%E2%80%93exclusion_principle


To see many other similar  (and different)  solved problems,  see the lessons

    - Counting elements in sub-sets of a given finite set
    - Advanced problems on counting elements in sub-sets of a given finite set
    - Challenging problems on counting elements in subsets of a given finite set
    - Selected problems on counting elements in subsets of a given finite set
    - Inclusion-Exclusion principle problems

in this site.


Happy learning (!)